# Strict Subgoal Execution: Reliable Long-Horizon Planning in Hierarchical Reinforcement Learning

### v1

@@ 4.2 Strict Subgoal Execution with Decoupled Exploration
$\pi^{h}(\tilde{g}_{t}|\tilde{s}_{t},g)=\begin{cases}\tilde{g}_{\max,t}:=\arg\max_{\tilde{g}\in\mathcal{G}}Q^{h}(\tilde{s}_{t},\tilde{g})&\text{with probability }1-\epsilon,\\ \text{randomly selected }\tilde{g}_{\mathrm{rand}}\in\mathcal{G}&\text{with probability }\epsilon,\end{cases}$ (2)
where $\tilde{g}_{\max,t}$ is the maximum subgoal, $\tilde{s}_{t}=(s_{t},t/H)$ augments $s_{t}$ with normalized time $t/H$ to reflect episode progress, and $Q^{h}$ is the high-level value function trained using RL. This formulation enables $\pi^{h}$ to explore the goal space extensively, independent of the agent’s current location.
Similar to prior work, we construct a waypoint path $(\mathrm{wp}_{1},\dots,\mathrm{wp}_{n})$ from the current state $s_{t}$ to the subgoal $\tilde{g}_{t}$ using the graph $G=(V,E)$. The low-level policy $\pi^{l}$ sequentially follows these waypoints with unlimited step execution until the subgoal is reached. Let $k_{t}$ denote the number of steps required to reach $\tilde{g}_{t}$ from $s_{t}$; if the agent fails, $k_{t}$ may extend up to the full episode length $H$. Based on this, the $i$-th high-level decision time step is defined as $t_{i}=t_{i-1}+k_{t_{i-1}}$ for $i\geq 1$ with $t_{0}=0$. Unlike previous methods that continue the episode regardless of subgoal completion, SSE terminates the episode if the subgoal is not successfully reached. A subgoal is considered reached if $||\phi(s_{t+k_{t}})-\tilde{g}_{t}||<\lambda$, where |||||$\cdot$||| | ⋅ | | denotes the Euclidean norm and $\lambda>0$ is a distance threshold. If this condition is met, the next subgoal is generated; otherwise, the episode ends with no reward. The corresponding high-level transition for training $\pi^{h}$ is:
$\begin{cases}(\tilde{s}_{t},g,\tilde{g}_{t},\sum_{j=t}^{t+k_{t}-1}r_{j},\tilde{s}_{t+k_{t}})&\text{if }||\phi(s_{t+k_{t}})-\tilde{g}_{t}||<\lambda,\\ (\tilde{s}_{t},g,\tilde{g}_{t},0,\tilde{s}_{T})&\text{otherwise},\end{cases}$ (3)
where $\tilde{s}_{T}=(s_{T},1)$ is the terminal state. Under this structure, only successful subgoal completions yield positive rewards, thereby encouraging reliable subgoal execution. To leverage partial successes, we adopt an idea inspired by hindsight experience replay (HER) [26]. If the agent fails to reach $\tilde{g}_{t}$ but reaches the final visited waypoint $\mathrm{wp}_{\mathrm{final}}$ in the waypoint path, we treat it as a virtual subgoal and record the following transition: $(\tilde{s}_{t},g,\mathrm{wp}_{\mathrm{final}},\sum_{j=t}^{t+k^{\prime}_{t}-1}r_{j},\tilde{s}_{t+k^{\prime}_{t}})$, where $k^{\prime}_{t}$ is the number of steps taken to reach $\mathrm{wp}_{\mathrm{final}}$, and it satisfies $||\phi(s_{t+k^{\prime}_{t}})-\mathrm{wp}_{\mathrm{final}}||<\lambda$. This allows the transition to be treated as successful and contributes positively to training. All successful and failed transitions are stored in a high-level replay buffer $\mathcal{B}^{h}$ and used to train $\pi^{h}$ and $Q^{h}$ via standard off-policy RL.
@@
While SSE encourages precise goal-directed behavior, its strict reward structure can limit exploration. To address this, we introduce a decoupled exploration policy $\pi^{\mathrm{exp}}$ that actively explores under-visited regions of the goal space. At each step $t$, $\pi^{\mathrm{exp}}$ selects a subgoal according to:
$\pi^{\mathrm{exp}}(\tilde{g}_{t}|\tilde{s}_{t},g)=\begin{cases}g&\text{with probability }\frac{1}{3},\\ \tilde{g}_{\max,t}&\text{with probability }\frac{1}{3},\\ \text{randomly selected }\tilde{g}_{\mathrm{novel}}\in C_{\mathcal{G}}^{m_{\mathrm{novel}}}&\text{with probability }\frac{1}{3},\end{cases}$ (4)
where $\tilde{g}_{\mathrm{novel}}$, referred to as the novel subgoal, is sampled from the least visited region $C_{\mathcal{G}}^{m_{\mathrm{novel}}}$ among $M$ grid cells $C_{\mathcal{G}}^{m}$ of size $d_{\mathcal{G}}$ that partition the goal space, $\mathcal{G}=\bigcup_{m=1}^{M}C_{\mathcal{G}}^{m}$. Here, $m_{\mathrm{novel}}$ is determined by $m_{\mathrm{novel}}=\arg\min_{m}N(C_{\mathcal{G}}^{m})$, where $N(C_{\mathcal{G}}^{m})$ denotes the number of episodes in which grid $C_{\mathcal{G}}^{m}$ has been visited. The policy $\pi^{\mathrm{exp}}$ follows the same SSE mechanism as $\pi^{h}$ and stores its transitions in the high-level buffer $\mathcal{B}^{h}$. This strategy promotes broad and effective exploration by encouraging visits to final goals, high-value subgoals, and novel subgoals in a balanced manner. Early in training, we rely solely on $\pi^{\mathrm{exp}}$ and gradually mix it with $\pi^{h}$ using a target ratio of $\eta:(1-\eta)$, where $\eta$ is the exploration control factor. By improving coverage of promising and under-explored regions without additional learning, $\pi^{\mathrm{exp}}$ enhances subgoal accessibility and high-level generalization.
To illustrate the behavior of the proposed method, Fig. 2 shows grid-wise visitation in the goal space and low-level trajectories toward various initial subgoals at three stages of training: (a) early (10K steps), (b) intermediate (150K steps), and (c) final (500K steps). The environment is a U-maze where a MuJoCo [43] Ant agent navigates to the final goal $g$ located at the upper left of the map. In (a), most of the goal space remains unexplored, and initial subgoals selected by $\pi^{\mathrm{exp}}$ and $\pi^{h}$, including the final goal, random subgoals, and novel regions, result in wide-ranging trajectories that promote broad exploration. As training progresses in (b), the agent expands its coverage across the map, and in (c), it consistently reaches the goal $g$. Notably, SSE enables the agent to reach any reachable subgoal in a single high-level step, regardless of distance, supporting reliable execution and efficient long-range planning. This allows the agent to solve complex tasks using only a few high-level decisions.
@@ 4.3 Failure-Aware Path Refinement
In the SSE framework, reliable subgoal execution is crucial since each subgoal must be reached within a single high-level step. As described in Section 2, graph-based methods use Dijkstra’s algorithm to compute waypoint paths from $\phi(s_{t})$ to $\tilde{g}_{t}$ on a graph $G=(V,E)$, with edge distances defined by $d$ in Eq. (1). However, this distance does not capture failure cases such as collisions or getting stuck, leading agents to often fail even along the shortest path. We observe that such failure-prone regions significantly hinder subgoal success. To address this, we introduce a failure-aware path refinement strategy that increases edge costs in unreliable regions, steering the planner toward safer alternatives.
Using the grid cells $C_{\mathcal{G}}^{m}$, we define $N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})$ as the number of episodes where the agent fails to reach a subgoal located outside the current cell, i.e., $\tilde{g}_{t}\in C_{\mathcal{G}}^{m^{\prime}}$ with $m^{\prime}\neq m$, but the episode terminates within $C_{\mathcal{G}}^{m}$. This indicates failure to exit the current region. To discourage such behavior, we increase traversal costs in these areas by defining the failure ratio as $\mathrm{ratio}_{\mathrm{fail}}(C_{\mathcal{G}}^{m})=N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})/N(C_{\mathcal{G}}^{m})$, and refine the edge distance from node $v_{1}$ to $v_{2}$ as:
@@
$\tilde{d}(v_{1}\rightarrow v_{2})=d(v_{1}\rightarrow v_{2})\times\max\left(1,c_{\mathrm{dist}}\cdot\mathrm{ratio}_{\mathrm{fail}}(C_{\mathcal{G}}^{m})\right),\quad\forall v_{2}\in C_{\mathcal{G}}^{m},\quad\forall m,$ (5)
where $d$ is the original edge distance defined in Eq. (1), and $c_{\mathrm{dist}}>1$ is a scaling factor. A higher failure ratio increases the adjusted distance, encouraging Dijkstra’s algorithm to avoid unreliable regions. If no alternative exists, the original path is still taken to preserve subgoal reachability.
@@
Fig. 3 illustrates the effect of the proposed path refinement in a bottleneck environment. In (a), without refinement, the agent repeatedly follows the shortest path through a narrow corridor near a wall, often resulting in failure. In (b), with refinement applied, increased edge costs in high-failure regions steer Dijkstra’s algorithm toward safer detours. When no alternatives exist (e.g., in the bottleneck), the agent still passes through, preserving reachability. This demonstrates that the refinement enhances subgoal success while maintaining overall reachability, supporting more stable execution in complex tasks. In summary, the proposed SSE framework is illustrated in Fig. 4, with a condensed version of the algorithm provided in Algorithm 1. The full algorithm and implementation details, including the graph construction method and training procedures, are provided in Appendix B.
@@
                   Sample a subgoal $\tilde{g}_{t}$ from $\pi^{h}$ or $\pi^{\mathrm{exp}}$ with ratio $\eta:(1{-}\eta)$
                  Plan the waypoint path from $\phi(s_{t})$ to $\tilde{g}_{t}$
                  Roll out low-level policy $\pi^{l}(s_{t},\mathrm{wp}_{i})$ $\forall i$
                  if $\|\phi(s_{t+k_{t}})-\tilde{g}_{t}\|<\lambda$ (success) then
                         Store $(\tilde{s}_{t},g,\tilde{g}_{t},\sum_{j=t}^{t+k_{t}-1}r_{j},\tilde{s}_{t+k_{t}})$
@@
                         Store $(\tilde{s}_{t},g,\tilde{g}_{t},0,\tilde{s}_{T})$ and terminate
@@ 5 Experiments
In this section, we evaluate the proposed SSE framework on 9 challenging long-horizon tasks, including 5 AntMaze environments (U-maze, $\pi$-maze, AntMazeComplex, AntMazeBottleneck, and AntMazeDoubleBottleneck). Among these, U-maze has a simple layout, AntMazeBottleneck features narrow corridors, and the others present more complex structures. We also assess 2 KeyChest tasks (AntKeyChest, AntDoubleKeyChest), where the agent must collect 1 or 2 keys before reaching the final goal, even though the keys are not explicitly defined as goals. Additionally, we evaluate 2 Reacher tasks (ReacherWall, ReacherWallDoubleGoal), where a 3D robot must navigate obstacles to reach one or two goals. KeyChest and DoubleGoal tasks require intermediate objectives, making them ideal for testing high-level planning. Fig. 5 provides visualizations, with further details presented in Appendix C. While the main experiments focus on fixed-goal settings, additional comparisons for random-goal setups are included in Appendix D.1.
@@ 5.1 Performance Comparison
We compare SSE with a range of baselines, including both hierarchical RL and recent graph-based methods. Specifically, we evaluate 2 HRL approaches: HIRO [40], which improves sample efficiency via hindsight goal relabeling, and HRAC [16], which penalizes subgoal selection based on reachability. We also include 3 graph-based HRL methods: HIGL [12], which applies intrinsic penalties over a graph, DHRL [19], which finds a path and gives waypoints to low-level as a goal via graph palnning, and NGTE [20], which expands graphs using novelty to better address fixed-goal settings. Additionally, we consider 2 graph-based methods without explicit high-level policies: PIG [42], which integrates graph structures into imitation learning to skip redundant subgoal actions, and BEAG [15], which uses breadth exploration with imaginary landmarks for goal-reaching. SSE is evaluated with the best hyperparameter settings ($c_{\mathrm{dist}}$, $d_{\mathcal{G}}$, $\eta$) from ablation studies, while all baselines use author-provided implementations. Detailed descriptions of each algorithm, along with the hyperparameter configurations for our proposed method, are provided in Appendix C.
@@
Fig. 6 shows success rate comparisons across all tasks, where solid lines indicate the mean over 5 seeds and shaded areas denote standard deviations. From the result, SSE consistently outperforms both graph-based and conventional HRL methods across all benchmarks. Conventional HRL methods (DHRL, HIRO, PIG, HRAC, and HIGL) rely on random goal sampling for exploration but often struggle with fixed-goal tasks, while occasionally succeeding in random-goal setups provided in Appendix D.1, highlighting the increased challenge of fixed goals due to limited exploration opportunities. In relatively simple environments such as U-maze, $\pi$-maze, and AntMazeComplex, baseline methods like BEAG and NGTE demonstrate reasonable performance, but SSE typically converges more quickly. In bottleneck environments, SSE further excels by using failure-aware path refinement to avoid unstable regions as shown in Fig. 3. In more complex tasks like KeyChest and ReacherWallDoubleGoal, which require reaching intermediate objectives, baseline methods largely fail. This is because conventional HRL suffers from long high-level horizons, and goal-centric methods without high-level decision-making, such as BEAG, cannot reason about intermediate targets. In contrast, SSE mitigates these issues by strictly enforcing subgoal completion, reducing high-level decision steps. These results highlight the robustness, efficiency, and generalization of the proposed SSE framework. For practical comparison, we also evaluate the computational complexity against major baselines in Appendix D.2. The results show that our method achieves lower complexity per iteration, demonstrating its superiority in terms of computational efficiency.

### v2

@@ 4.2 Strict Subgoal Execution with Frontier Experience Replay
$\pi^{h}(\tilde{g}_{t}\mid s_{t},g)=\begin{cases}\tilde{g}_{\max,t}:=\arg\max_{\tilde{g}\in\mathcal{G}}Q^{h}(s_{t},\tilde{g},g)&\text{with probability }1-\epsilon,\\ \tilde{g}_{\mathrm{rand}}\sim\mathrm{Uniform}(\mathcal{G})&\text{with probability }\epsilon,\end{cases}$ (3)
where $\tilde{g}_{\max,t}$ is the greedy subgoal, in practice generated by the actor network trained to choose the maximum of $Q^{h}$, $\tilde{g}_{\mathrm{rand}}$ is sampled uniformly from $\mathcal{G}$ to ensure persistent global exploration, and $Q^{h}$ is trained off-policy using $\mathcal{B}_{F}^{h}$. This formulation allows $\pi^{h}$ to select diverse points in the goal space independently of the agent’s current location, which is crucial for broader exploration.
Exploration Policy $\pi^{\mathrm{exp}}$: To promote exploration, $\pi^{\mathrm{exp}}$ targets low-density, underexplored regions of the goal space. To do this, we first identify novel regions $V_{\mathrm{novel}}\subset\mathcal{G}$, for which we employ either a grid-based estimator or a model-based approach. In this paper, we consider both: the grid-based estimator is used for simplicity in 2D/3D goal spaces that are widely adopted in real-world settings (e.g., position goal spaces), whereas the model-based approach, although more complex, is adopted to facilitate extension to higher-dimensional goal spaces. The exploration policy $\pi^{\mathrm{exp}}$ is defined as:
$\pi^{\mathrm{exp}}(\tilde{g}_{t}\mid s_{t},g)=\begin{cases}g&\text{with probability }\tfrac{1}{3},\\ \tilde{g}_{\max,t}&\text{with probability }\tfrac{1}{3},\\ \tilde{g}_{\mathrm{novel}}\sim\mathrm{Uniform}(V_{\mathrm{novel}})&\text{with probability }\tfrac{1}{3}.\end{cases}$ (4)
where the novel region $V_{\mathrm{novel}}$ is defined in one of two ways, depending on whether we use the grid-based or the model-based learning scheme, that is, $V_{\mathrm{novel}}=V_{\mathrm{novel}}^{\mathrm{grid}}$ or $V_{\mathrm{novel}}=V_{\mathrm{novel}}^{\mathrm{model}}$. The grid-based novel region is defined as $V_{\mathrm{novel}}^{\mathrm{grid}}:=C_{\mathcal{G}}^{m_{\mathrm{novel}}}$, where $C_{\mathcal{G}}^{m}$ denotes the grid cell of size $d_{\mathcal{G}}$ that partitions the goal space $\mathcal{G}$, $m$ indexes the grid cells, and $m_{\mathrm{novel}}=\arg\min_{m}N(C_{\mathcal{G}}^{m})$ is the index of the cell with the smallest visitation count $N(C_{\mathcal{G}}^{m})$. The model-based novel region is defined as $V_{\mathrm{novel}}^{\mathrm{model}}:=\left\{\arg\max_{v\in V}\lVert f_{\zeta}(v)-f_{\zeta_{\mathrm{targ}}}(v)\rVert\right\},$ where $f_{\zeta}$ is the learner network with parameter $\zeta$, and $f_{\zeta_{\mathrm{targ}}}$ is the initialized target network proposed by (Burda et al., 2018) to describe novel regions based on prediction error. The learner $f_{\theta}$ is trained to minimize the prediction error for node samples. Then, the novel subgoal $\tilde{g}_{\mathrm{novel}}$ is uniformly sampled from $V_{\mathrm{novel}}$. Both $\pi^{h}$ and $\pi^{\mathrm{exp}}$ operate under the same SSE mechanism, where episodes continue only upon successful subgoal completion and terminate on failure. To control exploration, we sample from $\pi^{\mathrm{exp}}$ early in training and then gradually mix it with $\pi^{h}$ using a ratio $\eta:(1-\eta)$, where $\eta$ controls exploration strength. This balanced scheme preserves coverage, accelerates the discovery of reachable subgoals, and improves high-level generalization. Implementation details and parameters such as $\eta$ are described in  B.2
@@
To illustrate the behavior of the proposed method, Fig. 2 shows grid-wise visitation in the goal space and low-level trajectories toward various initial subgoals at three stages of training: (a) early (10K steps), (b) intermediate (150K steps), and (c) final (500K steps). The environment is a U-maze where a MuJoCo (Todorov et al., 2012) Ant agent navigates to the final goal $g$ located at the upper left of the map. In (a), most of the goal space remains unexplored, and initial subgoals selected by $\pi^{\mathrm{exp}}$ (grid-based) and $\pi^{h}$, including the final goal, random subgoals, and novel regions, result in wide-ranging trajectories that promote broad exploration. As training progresses in (b), the agent expands its coverage across the map, and in (c), it consistently reaches the goal $g$. Notably, SSE enables the agent to reach any reachable subgoal in a single high-level step, regardless of distance, supporting reliable execution and efficient long-range planning. This allows the agent to solve complex tasks using only a few high-level decisions.
@@ 4.3 Failure-Aware Path Refinement
In the SSE framework, reliable subgoal execution is crucial as each subgoal must be reached within a single high-level step. As described in Section 2, graph-based methods use Dijkstra’s algorithm to compute waypoint paths from $\phi(s_{t})$ to $\tilde{g}_{t}$ on a graph $G=(V,E)$, with edge distances defined by $d$ in equation 1. However, this distance ignores failure cases like collisions or getting stuck, causing agents to fail even on the shortest path. We observe these failure-prone regions significantly hinder subgoal success. To address this, we introduce a failure-aware path refinement strategy that increases edge costs in unreliable regions, steering the planner toward safer alternatives. For each edge from node $v_{1}$ to node $v_{2}$, we define the failure ratio $\mathrm{ratio}_{\mathrm{fail}}$ for the target node $v_{2}$, and refine the edge by increasing its distance as the failure ratio increases as
@@
$\tilde{d}(v_{1}\rightarrow v_{2})=d(v_{1}\rightarrow v_{2})\times\max\left(1,c_{\mathrm{dist}}\cdot\mathrm{ratio}_{\mathrm{fail}}(v_{2})\right),~\forall v_{1},~v_{2}\in V,$ (5)
where $d$ is the original edge distance from equation 1, and $c_{\mathrm{dist}}>1$ is a scaling factor. For consistency with our exploration policy, the failure ratio for the target node is also defined in two ways, following the grid-based and model-based approaches: $\mathrm{ratio}_{\mathrm{fail}}=\mathrm{ratio}_{\mathrm{fail}}^{\mathrm{grid}}$ or $\mathrm{ratio}_{\mathrm{fail}}=\mathrm{ratio}_{\mathrm{fail}}^{\mathrm{model}}$. The grid-based failure ratio is defined as $\mathrm{ratio}_{\mathrm{fail}}^{\mathrm{grid}}(v_{2}):=N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})/N(C_{\mathcal{G}}^{m}),$ where $v_{2}\in C_{\mathcal{G}}^{m}$, $N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})$ is the number of failure episodes in the cell $C_{\mathcal{G}}^{m}$. The model-based failure ratio is defined as $\mathrm{ratio}_{\mathrm{fail}}^{\mathrm{model}}(v_{2}):=F_{\xi}(v_{2}),$ where $F_{\xi}$ is a failure predictor network with parameters $\xi$ trained over node samples using a cross-entropy objective, assigning label $1$ to nodes from failed trajectories and label 0 to nodes from successful trajectories. To ensure sufficient low-level policy competence, the failure-aware path refinement is activated only after $\lambda_{\mathrm{count}}$ successful visits. A higher failure ratio increases the adjusted edge distance, encouraging Dijkstra’s algorithm to avoid unreliable regions.
As described above, we instantiate SSE with two variants that differ in their exploration policy and failure-aware path refinement, namely a grid-based approach and a model-based approach. We refer to SSE with the grid-based variant as SSE (Grid) and SSE with the model-based variant as SSE (Model), and we empirically compare these two variants to analyze how the choice of approach affects learning performance.
@@
Fig. 3 illustrates the effect of the proposed grid-based path refinement in a bottleneck environment. In (a), without refinement, the agent repeatedly follows the shortest path through a narrow corridor near a wall, often resulting in failure. In (b), with refinement applied, increased edge costs in high-failure regions steer Dijkstra’s algorithm toward safer detours. When no alternatives exist (e.g., in the bottleneck), the agent still passes through, preserving reachability. This demonstrates that the refinement enhances subgoal success while maintaining overall reachability, supporting more stable execution in complex tasks. In summary, the proposed SSE framework is illustrated in Fig. 4, with a condensed version of the algorithm provided in Algorithm 1. The full algorithm and implementation details, including the graph construction and training losses for RL training and model-based approach are provided in Appendix B.
@@
         Sample a subgoal $\tilde{g}_{t}$ from $\pi^{h}$ and $\pi^{\mathrm{exp}}$
        Plan a waypoint path from $\phi(s_{t})$ to $\tilde{g}_{t}$
        Roll out low-level policy $\pi^{l}(s_{t},\mathrm{wp}_{i})$ $\forall i$:
       if $\|\phi(s_{t^{\prime}})-\tilde{g}_{t}\|<\lambda$ (success) then
            Store the success transition and continue
@@
            Store the failure and partial-success transitions, then terminate the episode
@@
  Count $N(C_{\mathcal{G}}^{m})$ and $N_{\mathrm{fail}}(C_{\mathcal{G}}^{m})~\forall m$ (Grid-based)
  Update models $f_{\zeta}$ and $F_{\xi}$ (Model-based)
@@ 5 Experiments
In this section, we evaluate our SSE framework on 9 challenging long-horizon tasks, including 5 AntMaze environments (U-maze, $\pi$-maze, AntMazeComplex, AntMazeBottleneck, and AntMazeDoubleBottleneck). These range from simple layouts (U-maze) to complex structures with narrow corridors (AntMazeBottleneck). We also assess 2 KeyChest tasks (AntKeyChest, AntDoubleKeyChest), where the agent must collect 1 or 2 keys before reaching the final goal, even though the keys are not explicitly defined as goals. Additionally, we evaluate 2 Reacher tasks (ReacherWall, ReacherWallDoubleGoal), where a 3D robot must navigate obstacles to reach one or two goals. KeyChest and DoubleGoal tasks require intermediate objectives, making them ideal for testing high-level planning. See Fig. 5 for visualizations and Appendix C for details. While the main experiments focus on fixed-goal settings, additional comparisons for random-goal setups are included in Appendix D.1.
@@ 5.1 Performance Comparison
We compare SSE with a range of hierarchical RL and recent graph-based methods. Specifically, we evaluate 2 HRL approaches: HIRO (Nachum et al., 2018b), which improves sample efficiency via hindsight goal relabeling, and HRAC (Zhang et al., 2020), which penalizes subgoal selection based on reachability. We also include 3 graph-based HRL methods: HIGL (Kim et al., 2021), which applies intrinsic penalties over a graph, DHRL (Lee et al., 2022), which finds a path and gives waypoints to low-level as a goal via graph planning, and NGTE (Park et al., 2024), which expands graphs using novelty to better address fixed-goal settings. Additionally, we consider 2 graph-based methods without explicit high-level policies: PIG (Kim et al., 2023), which integrates graph structures into imitation learning to skip redundant subgoal actions, and BEAG (Yoon et al., 2024), which uses breadth exploration with imaginary landmarks for goal-reaching.
For our proposed framework, we evaluate two variants: a grid-based implementation SSE (Grid), which offers a simple implementation in practical 2D/3D goal spaces, and a learning-based implementation SSE (Model), which is designed to be scalable to higher-dimensional goal spaces, as described in Section 4. Both variants are evaluated with the best hyperparameter settings ($c_{\mathrm{dist}}$, $d_{\mathcal{G}}$, $\eta$) from ablation studies, while all baselines use author-provided implementations. In FER, the threshold $\lambda$ used to determine subgoal reachability is the same as the threshold used to define the low-level reward in Section 2.2. We use the same $\lambda$ values as the baseline methods to ensure a fair and consistent comparison. Detailed descriptions of each algorithm, along with the hyperparameter configurations for our proposed method, are provided in Appendix C.
@@
Fig. 6 shows mean success rates over 5 seeds (solid lines) with standard deviations (shaded). From the results, both SSE variants consistently outperform other graph-based and conventional HRL methods across all benchmarks. Conventional HRL methods (DHRL, HIRO, PIG, HRAC, and HIGL) rely on random goal sampling for exploration but often struggle with fixed-goal tasks, while occasionally succeeding in random-goal setups provided in Appendix D.1, highlighting the increased challenge of fixed goals due to limited exploration opportunities. In relatively simple environments such as U-maze, $\pi$-maze, and AntMazeComplex, baseline methods like BEAG, grid-based nodes, and NGTE demonstrate reasonable performance, but SSE typically converges more quickly. In bottleneck environments, SSE further excels by using failure-aware path refinement to avoid unstable regions as shown in Fig. 3. In more complex tasks like KeyChest and ReacherWallDoubleGoal, which require reaching intermediate objectives, baseline methods largely fail. Conventional HRL suffers from long high-level horizons, and goal-centric methods without high-level decision-making, such as BEAG, cannot reason about intermediate targets. In contrast, SSE mitigates these issues by strictly enforcing subgoal completion, reducing high-level decision steps. These results highlight the versatility, efficiency, and generalization of the proposed SSE framework. While both variants succeed in all tasks, SSE (Grid) converges faster due to lower overhead. In contrast, SSE (Model) is designed for scalability to high-dimensional spaces where discretization is infeasible, despite higher training complexity. In addition, extended evaluations on AntPush, AntFall, and Ant4Rooms in Appendix E.3 confirm that SSE consistently matches or exceeds baselines. Furthermore, computational analysis in Appendix D.2 demonstrates that SSE achieves lower complexity per iteration than baselines.